Suppose that you are computing an RSA key pair.   (Marks 6)        What are ? and ?...

  1. Suppose that you are computing an RSA key pair.   (Marks 6)        What are ? and ? and ?(?) for an ? = 51?

       Find a legal RSA public key pair for this ? and ?.

       How many possible values for ?? are there?

In a RSA cryptosystem, a person Renee uses two prime numbers p = 13        and q = 17 to generate her public and private keys. If the public key of       Renee is 35, then the private key of Renee is _______.    
Answer #1
  • Given the prime numbers p and q, and the public key e = 35, The private key can be obtained by finding the multiplicative inverse of e mod ((p-1)*(q-1)).
  • The multiplicate inverse is a number d, such that d<(p-1)(q-1) and ed mod (p-1)(q-1) = 1 mod (p-1)(q-1).
  • here e = 35, p = 13 , q = 17.
    • 35*d mod (12*16) = 35*d mod*192 = 1 mod 192.
    • The value of d has to be checked one by one from 1 to 191. This long computation process is exactly why RSA is efficient.
    • For d = 11 we get 35*11 mod 192 = 385 mod 192 = 1 = 1 mod n
  • 11, is the multiplcative inverse of 35 mod 192 and it is the private key for the given system